# LLM-Enabled In-Context Learning for Data Collection Scheduling in UAV-assisted Sensor Networks

### v1

@@ V Attack Model
This section presents the defined attack model and procedure. In particular, the attack model targets LLM-enabled ICL used for UASNETs. The attacker uses black-box jailbreaking, specifically the exploitation of template completion attacks and context-based manipulation, to exploit vulnerabilities in the edge-based LLM that is responsible for optimizing data collection scheduling. The attacker has no access to the internal model parameters, gradients or weights, but can query the system via external API interactions. The attacker manipulates the LLM’s data collection scheduling prompts to reduce the performance of the system. The attacker interacts with the LLM via a standard query interface. In this approach, the attacker crafts a malicious sample with context-based manipulation. This technique exploits the ICL by tricking the LLM into adopting malicious behavior in data collection scheduling through manipulated examples. The attacker can bias the UAV data collection schedule, prioritize compromised ground sensors, or degrade packet loss through context-based manipulation. In the following, we present the attack procedure step by step:
@@
    When the LLM processes the malicious prompt, it generates a biased schedule: ”UAV will visit Sensor Z (bad Channel) first, then Sensor X”. This output reflects the attacker’s influence, as Sensor Z is prioritized over others, despite its bad channel. The attacker’s manipulation disrupts the system normal operation, leading to degraded data collection performance.
The above attack process demonstrates how an adversary can exploit the LLM’s reliance on ICL to manipulate its behavior without needing access to its internal parameters or weights [29]. Table I provides a summary of the attack model.
@@ VI Proposed ICL-Based Data Collection Scheduling
Input: UAV sensory data $S={s_{j}(q_{j},\gamma_{j}),j=1..N}$, edge-hosted LLM
@@
UAV collects sensory data $\mathcal{S}$ with queues length and channels condition;;
@@
Otherwise, UAV sends sensory data $\mathcal{S}$ to the LLM for decision-making (exploitation);
@@
Adversary attempts to modify LLM’s task description.;
@@
Despite the context-based manipulation attempt, the LLM analyzes the collected sensory data $\mathcal{S}$ according to its trained model.;
@@
Ensure scheduling adherence, preventing the model from being coerced into unrelated roles;
@@
Update feedback based on environment state, past decisions, and system performance;
return Optimized data collection schedule;
Algorithm 1 Proposed ICLDC with Epsilon-Decay
This section introduces the proposed ICL-based Data Collection Scheduling (ICLDC) scheme.
ICL enables LLMs to enhance their performance on specific tasks by leveraging structured natural language inputs, such as task descriptions and solution demonstrations. This process can be formally represented by Eq. 5 where $TD_{\text{task}}$ denotes the task description, $Ex_{t}$ represents the set of examples at time $t$, $e_{t}$ corresponds to the environmental state associated with the target task at time $t$, LLM refers to the large language model, and $a_{t}$ denotes the model output.
@@
For sequential decision-making problems, the LLM is expected to process the initial task description $TD_{\text{task}}$, extract feedback from the example set $Ex_{t}$, and generate decisions $a_{t}$ based on the current environment state $e_{t}$.
Table III: Configuration
Parameters Values
Number of ground sensors ($N$) 10
Queue length ($D$) 40
Number of Time Steps 20
Epsilon decay ($\epsilon$) 0.398
Maximum transmit power 100 mW
maximum battery capacity 50 J
$TD_{\text{task}}$ is crucial to provide target task information to the LLM model. In particular, it involves task goals, input data, rules and constraints, output, and feedback loop. This task description avoids the complexity of dedicated optimization model design and relieves the operator from expert knowledge on optimization techniques.
Figure 3: Performance analysis of ICLDC in different scenarios with 10 ground sensors concerning the number of lost packets per time step. (a) The network cost at each time step of ICLDC and maximum channel gain, (b) The network cost at each episode of ICLDC with different LLMs (c) The network cost at each time step of ICLDC and with different epsilon-decay values, (d) Normal and Under attack operation of ICLDC where the jailbreaking attack happens.
VI-A Proposed ICLDC
Figure 4: Performance with changing the number of ground sensors.
Algorithm 1 depicts the proposed ICLDC where ICL is used to schedule data collection. Detailed explanations are provided as follows:
@@ VI-A Proposed ICLDC
    Data Collection: The UAV collects sensory data, including queue length, and channel condition[30],[31], and employs an epsilon decay parameter ($\epsilon$) to balance exploration and exploitation in the submission of collected data to the edge-hosted LLM.
@@
    Context-based Manipulation Attack: The adversary attempts to modify the LLM’s task description by injecting malicious context or altering its instructions to mislead the LLM into making suboptimal decisions. This manipulation seeks to deceive the LLM into activating incorrect sensors. The attacker could modify the task description to give higher priority to a sensor with poor signal strength, leading to inefficient data collection and transmission.
@@
    Contextual Understanding: Despite the context-based manipulation attempt, the LLM analyzes the collected sensory data according to its trained model.
@@
    Data Collection Scheduling: The edge LLM uses ICL to analyze the gathered data and schedule data collection that minimizes packet loss. The LLM is restricted to data collection scheduling tasks and cannot be coerced into adopting unrelated or harmful roles.
@@
    Adaptive Learning and System Update: The LLM continuously updates its data collection scheduling based on new data received from the ground sensors. The LLM records its decisions, environment state, and system performance to derive feedback. When the next environment state arises, relevant feedback is retrieved to guide the LLM in making subsequent decisions.
@@ VII-A Implementation of ICLDC
We consider that N ground sensors are randomly deployed in a 100m × 100m area. The simulation is conducted in a custom Python 3.5 environment. The UAV is simulated and its actions are controlled by policies learned through ICLDC. Each ground sensor is modeled with a maximum battery capacity of 50 J and a maximum transmit power of 100 mW. Wireless communication between the UAV and sensors is simulated using the probabilistic channel model explained in Section IV. All experiments are run on a Lenovo workstation with Ubuntu 20.04 LTS, an Intel Core i5-7200U CPU @ 2.50GHz × 4, and 16 GIB RAM.
@@ VII-B Baselines
Maximum Channel Gain: The UAV collects logged sensor data from sensor nodes on the ground and transmits this contextual information to the LLM. The LLM uses this information to create a scheduling strategy using ICL that favors the sensor with the highest channel gain and not consider the queue length of ground sensors. This approach minimizes packet loss by tracking reliable communication links.

### v2

@@ V Attack Model
This section presents the defined attack model and procedure. In particular, the attack model targets LLM-assisted ICL used for UASNETs. The attacker uses black box jailbreaking, specifically template completion attacks and context-based manipulation to exploit vulnerabilities in the edge-based LLM responsible for optimizing the data collection schedule.
The attacker operates under a black box threat model and has no access to the internal model parameters, gradients, or weights, but can freely query the system via its external API interface. The overall goal of the attacker is to degrade system performance by manipulating the LLM’s requests for data collection and thus indirectly maximizing a network loss function $L$ (e.g., packet loss). Formally, let $S=f(\text{prompt})$ be the schedule generated by the LLM for a given prompt. The attacker crafts a malicious demonstration $\mathcal{M}$ to achieve the goal expressed in Eq. 6 subject to the constraint that only the prompt context is modified.
$\max_{\mathcal{M}}\;L\bigl(f(\text{prompt}\cup\mathcal{M})\bigr)$ (6)
For instance, a feasible instantiation of $\mathcal{M}$ (e.g., “prioritize sensors with bad channel conditions”) can coerce the LLM into a schedule $S$ that selects the worst-quality communication link, thereby substantially increasing $L$.
From a probabilistic perspective, let $O$ denote the LLM’s stochastic output model with unknown parameters. Given a benign prompt, the model produces a distribution over possible scheduling actions $y$. By appending the malicious demonstration $\mathcal{M}$, the attacker seeks to increase the conditional probability of a harmful action as expressed in Eq. 7 relative to the benign case without $\mathcal{M}$.
$P\bigl(y=\text{``choose low-gain sensor first''}\;\big|\;\text{prompt}\cup\mathcal{M}\bigr)$ (7)
The core heuristic in black-box ICL attacks is that the demonstration $\mathcal{M}$ induces a posterior shift in the model’s preference, effectively directing the model to different outputs that match the malicious example (Eq.8).
$\arg\max_{y}P(y\mid\text{prompt}\cup\mathcal{M})\;\;\neq\;\;\arg\max_{y}P(y\mid\text{prompt})$ (8)
This mechanism summarizes the essence of context-based manipulation: the injected $\mathcal{M}$ changes the effective decision boundary of the LLM without changing its internal parameters.
In practice, this technique exploits the ICL paradigm by tricking the LLM into adopting malicious scheduling behavior through manipulated in-context examples. In this way, the attacker can distort the data collection schedule of the drones, force the prioritization of compromised ground sensors, or systematically degrade packet loss, ultimately undermining the operational efficiency of UASNETs.
Figure 3: Closed-loop workflow for the proposed ICLDC. The system integrates an edge server-based LLM module with a safety verifier and an attack detection module. The process begins with the UAV collecting sensory data (e.g., queue length, channel state, battery level) and feedback. This data structures a task description for the LLM, which proposes an optimized schedule via ICL. The safety checker validates this schedule against predefined rules, overriding it if necessary. Concurrently, the attack detection module estimates jailbreaking probability via prompt perplexity analysis, triggering a random safe action upon suspicion. Executed schedules and performance metrics are logged into a feedback loop, enabling continuous LLM adaptation to minimize packet loss.
The attack procedure step by step is as follows:
@@
    When the LLM processes the malicious prompt, it generates a biased schedule: ”UAV will visit Sensor Z (bad Channel) first, then Sensor X”. This output reflects the attacker’s influence, as Sensor Z is prioritized over others, despite its bad channel. The attacker’s manipulation disrupts the system’s normal operation, leading to degraded data collection performance.
The above attack process demonstrates how an adversary can exploit the LLM’s reliance on ICL to manipulate its behavior without needing access to its internal parameters or weights [39]. Table III provides a summary of the attack model. Particularly, black box attacks represent a well-established and wide-ranging class of threats, as documented in numerous comprehensive surveys and studies [40], [41]. These attacks cover a wide range of techniques, including query, decision, and transfer-based methods, and are particularly relevant in scenarios involving closed-source models such as GPT-4o. We deliberately chose to study black-box attacks because working with proprietary models such as GPT-4o, whose internal parameters, architectures, and training data are not publicly available, comes with real limitations. Under such conditions, white-box attacks, which require access to model weights or gradients, are inherently unfeasible.
@@ VI Proposed ICL-Based Data Collection Scheduling
Input: UAV sensory data $\{e_{j}(q_{j},\gamma_{j},b_{j}),j=1..N\}$, edge-hosted LLM
@@
UAV collects sensory data $\{e_{j}\}$ with queues length, battery level and channels condition;
Compute perplexity $\mathcal{P}$ and attack likelihood $P_{attack}$;
Compute $\epsilon\leftarrow\epsilon+(1-\epsilon)*P_{attack}$;
@@
Otherwise, UAV sends sensory data $\{e_{j}\forall_{j}\}$ to the LLM for decision-making (exploitation).
@@
The adversary attempts to modify LLM’s task description.
@@
Despite the context-based manipulation attempt, the LLM analyzes the collected sensory data $\{e_{j}\forall_{j}\}$ according to its trained model.
@@
Ensures scheduling adherence, preventing the model from being coerced into unrelated roles;
A verifier ensures safe UAV operations by scoring and overriding LLM-generated schedules based on safety rules.
@@
Update feedback based on environment state, past decisions, and system performance
return optimized data collection schedule.
Algorithm 1 Proposed ICLDC with Perplexity-based Exploration Strategy
This section introduces the proposed ICL-based Data Collection Scheduling (ICLDC) scheme. We developed LLM prompts for decision-making using ICLDC by first defining clear goals that align with mission-critical tasks, such as the data collection schedule. Contextual information and domain-specific examples were embedded in the prompts to improve the LLM’s understanding of UAV scenarios and increase precision in complex tasks. Through iterative refinement based on model feedback and simulations, the prompts were refined to reduce ambiguity and better adapt to the UAV environment. To maintain continuity in dynamic missions, multi-turn interaction capabilities were integrated, while real-time responsiveness was evaluated to ensure low-latency, context-aware outputs suitable for edge deployment. Together, these steps have improved LLM’s effectiveness in supporting UASNETs. Our solution for managing multi-turn interactions involves the use of feedback loops. We incorporate feedback into the LLM at each round of conversation to iteratively refine the prompts. Continuously improving the prompts based on this feedback enhances the LLM’s ability to generate coherent and relevant responses.
In general, ICL enables LLMs to enhance their performance on specific tasks by leveraging structured natural language inputs, such as task descriptions and solution demonstrations. This process can be formally represented by Eq. 9 where $TD_{\text{task}}$ denotes the task description, $Ex_{t}$ represents the set of examples at time $t$, $e_{t}$ corresponds to the environmental state associated with the target task at time $t$, LLM refers to the large language model, and $a_{t}$ denotes the model output.
@@
For sequential decision-making problems, the LLM is expected to process the initial task description $TD_{\text{task}}$, extract feedback from the example set $Ex_{t}$, and generate decisions $a_{t}$ based on the current environment state $e_{t}$. $TD_{\text{task}}$ is crucial to provide target task information to the LLM model. In particular, it involves task goals, input data, rules and constraints, output, and a feedback loop. This task description avoids the complexity of dedicated optimization model design and relieves the operator from expert knowledge on optimization techniques.
However, as referred before, the reliance on contextual prompts introduces a vulnerability to context-based manipulation attacks. A malicious actor may inject a poisoned demonstration $\mathcal{M}$ into the example set $Ex_{t}$, effectively modifying the input as in Eq. 10.
$\text{LLM}:\quad TD_{\text{task}}\times(Ex_{t}\cup\mathcal{M}_{\text{mal}})\times e_{t}\rightarrow a_{t}^{\prime}$ (10)
The adversarial objective is to shift the LLM’s output distribution, steering the decision $a_{t}^{\prime}$ in Eq. 11 toward a suboptimal or harmful action (e.g., selecting a sensor with a critically low battery or poor channel gain).
$a_{t}^{\prime}$ $=\arg\max_{y}P(y\mid(TD_{\text{task}},Ex_{t}\cup\mathcal{M},e_{t}))$ (11)
$\neq a_{t}=\arg\max_{y}P(y=a_{t}\mid(TD_{\text{task}},Ex_{t},e_{t}))$
In ICLDC we mitigate the attack impact incorporating a safety verifier module that operates as a deterministic function $V(a_{t},e_{t})\rightarrow a_{t}^{\text{ver}}$. This module scores the proposed action $a_{t}$ against predefined safety rules. If $\text{Score}(a_{t})$ falls below a safety threshold $\tau_{safe}$ the verifier overrides the LLM’s decision and selects a sensor that maximizes $Score()$ (Eq. 12). This ensures that only safe actions are executed, thereby preserving system integrity even under adversarial prompting.
$a_{t}^{\text{ver}}=\arg\max_{j}\text{Score}(j)$ (12)
Moreover, an Attack Detection Module computes the perplexity $\mathcal{P}$ of the incoming prompt context $(TD_{\text{task}}\times Ex_{t})$ to identify potential manipulations. A high perplexity value $\mathcal{P}>\tau_{perplex}$ triggers an increase in the exploration parameter $\epsilon$ of an $\epsilon$-greedy policy, increasing the probability of taking a random safe action instead of following the potentially compromised LLM output.
VI-A ICLDC Workflow
Fig. 3 shows the closed-loop interaction between the UAV, the edge server hosting an LLM, and the ground sensors. The process starts with the data collection phase, where the UAV collects logged sensory data (e.g., queue length, channel condition, battery level) from the ground sensors. This data, along with feedback, is used to create a structured task description for the LLM. The LLM module processes this prompt through ICL to optimize the data collection schedule. The optimized data collection schedule is then evaluated by the safety checker, which matches it against predefined safety rules. If the schedule is deemed unsafe, the verifier overrides it and selects a safe alternative, resulting in a verified schedule. This verified command is sent back to the drone for execution.
At the same time, an attack detection module monitors the prompt. It calculates the perplexity of the prompt to estimate the probability of a jailbreaking attack. In case of a suspected attack, it adjusts the exploration parameter $\epsilon$ and causes the system to prioritize a random safe action over the potentially compromised output of the LLM. Finally, the system’s performance and the resulting state after executing the schedule are logged to generate feedback. This feedback allows the LLM to learn from past decisions and improve future data collection schedules, minimizing packet loss.
Algorithm 1 depicts the proposed ICLDC, where ICL is used to schedule data collection. Detailed explanations are provided as follows:
@@ VI-A ICLDC Workflow
    Data Collection: The UAV collects logged sensor data, including battery level, queue length, and channel condition. This data is transmitted to the edge server. The system uses a perplexity-based exploration strategy to balance exploration and exploitation before transmitting the data to the LLM. The perplexity of the prompt is calculated to assess its unusualness; a higher value may indicate a potential attack. An attack probability is then calculated based on this perplexity. Consequently, the epsilon value is dynamically adjusted in the epsilon greedy algorithm. If an attack is suspected, epsilon increases so that the system is more likely to perform a random safe action (exploration) instead of trusting a potentially malicious LLM output (exploitation).
@@
    Context-based Manipulation Attack: The adversary executes a black-box jailbreaking attack by manipulating the task description. This is achieved through context-based manipulation, a technique where a malicious demonstration $\mathcal{M}$ is injected into the prompt to mislead the LLM into generating a biased and suboptimal data collection schedule. The attacker’s objective is to indirectly maximize the network loss function $L$ (e.g., packet loss) by solving the optimization problem:
    $\max_{\mathcal{M}}\;L\left(f(\text{prompt}\cup\mathcal{M})\right)$ (13)
    where $f$ represents the LLM’s scheduling function. For instance, a malicious demonstration $\mathcal{M}$ such as “prioritize sensors with the worst channel conditions” conditions the LLM’s output distribution, increasing the probability:
    $P\left(y=\text{``choose low-gain sensor first''}\;\big|\;\text{prompt}\cup\mathcal{M}\right)$ (14)
    This posterior shift, formalized as $\arg\max_{y}P(y\mid\text{prompt}\cup\mathcal{M})\neq\arg\max_{y}P(y\mid\text{prompt})$, steers the model to prioritize compromised sensors with poor channel quality. The resulting schedule induces inefficient data collection and transmission, leading to significantly increased packet loss and degradation of overall network performance.
@@
    Contextual Understanding and Scheduling: The LLM (e.g., GPT-4o-mini) deployed at the edge server processes structured queries transmitted by the UAV and composed of $TD_{\text{task}}$, an exemplary set $Ex_{t}$ and the currently logged sensory data vector $e_{t}=\{q_{j}(t),\gamma_{j},b_{j}(t)\}$ for all $j$ sensors. The model leverages its inherent ICL capability to interpret this combined input.
    Despite potential context-based manipulation attacks, where a malicious demonstration $\mathcal{M}$ may be injected into the prompt to form $Ex_{t}\cup\mathcal{M}$, the LLM analyzes the entire context, including any malicious examples, to generate a probability distribution over all possible scheduling actions $y$. The core output is a data collection schedule $a_{t}$, which is the action with the highest conditional probability:
    $a_{t}=\arg\max_{y}P(y\mid(TD_{\text{task}},Ex_{t},e_{t})).$ (15)
    However, under attack, the adversarial context $\mathcal{M}$ induces a posterior shift, aiming to make:
    $a_{t}^{\prime}=\arg\max_{y}P(y\mid(TD_{\text{task}},Ex_{t}\cup\mathcal{M},e_{t}))$ (16)
    a harmful action, such as selecting a sensor with a critically low channel gain $\gamma_{j}<\gamma_{\text{th}}$.
    This proposed schedule $a_{t}$ (or $a_{t}^{\prime}$ if under attack) is not executed directly but is instead passed to the safety verifier for rigorous validation against predefined safety rules, ensuring system integrity is maintained before any action is taken.
@@
    Safety Verification and Override: The optimized data collection schedule is evaluated by the safety verifier. This module uses a scoring system based on critical safety rules (e.g., queue length, channel condition, battery level) to evaluate the selection of the LLM. If the schedule violates safety principles, i.e., receives a low safety score, the verifier overrides the decision. It then selects the sensor with the highest safety rating and creates a verified schedule that ensures safe and logical actions. The verifier and the LLM fulfill complementary tasks: The verifier enforces safety through predefined rules, while the LLM provides the adaptability and reasoning for complex, dynamic decisions.
@@
    Execution and Adaptive Learning: The verified schedule output by the safety verifier is dispatched to the UAV to execute the schedule. Following execution, the system’s performance metrics, specifically the packet loss resulting from the action, and the new environment state (e.g., updated queue lengths, channel conditions, and battery levels of all ground sensors) are recorded. When a subsequent scheduling decision is required at the next time step, the system constructs a new prompt. This prompt ingeniously combines the current sensory data with relevant feedback. The LLM leverages these in-context examples to refine its reasoning, avoid repeating past mistakes, and improve the quality of future data collection schedules, thereby guiding subsequent decisions to minimize packet loss.
The proposed ICLDC primarily involves text input, which significantly reduces the communication effort compared to multimodal settings (e.g., audio or visual-based queries). To improve communication efficiency, input token reduction techniques can be used, where unimportant or redundant tokens in the input sequence are removed before being passed to the LLM for inference. This form of text token pruning preserves the quality of inference while significantly reducing the amount of data to be transmitted. This directly results in lower energy costs for transmission for UAVs and reduces the impact on battery life and mission duration[42].
The extension of the proposed ICLDC to multi-UAV-assisted sensor networks presents a scalable solution for optimizing collaborative data collection in dynamic environments. Unlike conventional methods requiring complex training, ICL facilitates real-time decision-making by interpreting natural language inputs about network states. This approach addresses critical swarm challenges, including power control and velocity control, while bridging the simulation-to-reality gap through its inherent adaptability to unseen scenarios. However, practical deployment requires ensuring robust inter-UAV coordination.
Prompt optimization is a central and intentional part of the methodology, with prompts treated as fixed, engineered components akin to trained model parameters. Once optimized, they function as stable in-context policies, making concerns about prompt fragility less relevant. This design ensures consistent, reliable, and repeatable system performance across tasks.
While LLM inference introduces additional energy costs, our approach strategically mitigates this through prompt optimization, text-based inputs, and careful system design. By selecting efficient models, leveraging quantization, and optimizing batch sizes, inference overhead can be significantly reduced. These measures, combined with the lower resource demands of text over multimodal data, make the trade-off favorable for energy-efficient UASNETs in real-world deployments.
@@ VII Numerical Results and Discussions
Table IV: Simulation parameter settings
Parameters Values Units
Number of ground sensors ($N$) 10 [#]
Queue length ($D$) 40 [#]
Number of Time Steps 20 [#]
Maximum transmit power 100 [mW]
maximum battery capacity 50 [J]
Figure 4: Performance analysis of ICLDC in different scenarios with 10 ground sensors concerning the number of Packet Loss (a) The network cost at each time step of ICLDC and baselines: DQN and maximum channel gain, (b) The network cost at each time step of ICLDC with different LLMs (c) Performance with changing the number of ground sensors. (d) Normal and under attack operation of ICLDC, where the jailbreaking attack happens along with MQNS.
Figure 5: Measured LLM inference latency heatmap superimposed on UAV circular flight path
@@ VII-A Implementation of ICLDC
The key simulation parameters are shown in Table IV. We consider that 10 ground sensors are randomly deployed in a 100m × 100m area. Each ground sensor is modeled with a maximum battery capacity of 50J, a maximum transmission queue length of 40, and a maximum transmit power of 100mW. For the learning settings, we consider 20 time steps in each episode. The simulation is conducted in a custom Python 3.5 environment. The UAV is simulated and its actions are controlled by policies learned through ICLDC. Wireless communication between the UAV and sensors is simulated using the probabilistic channel model explained in Section IV. All experiments are run on a Lenovo workstation with Ubuntu 20.04 LTS, an Intel Core i5-7200U CPU @ 2.50GHz × 4, and 16 GB RAM.
@@ VII-B Baselines
Maximum Channel Gain: The UAV collects logged sensor data from sensor nodes on the ground and transmits this contextual information to the LLM. The LLM uses this information to create a scheduling strategy using ICL that favors the sensor with the highest channel gain and does not consider the queue length of ground sensors. This approach minimizes packet loss by tracking reliable communication links. Max-Queue-Neighbor-Selection (MQNS): MQNS makes use of spatial awareness by restricting decision-making to a predefined neighborhood around the UAV’s current location. The neighborhood is defined by a configurable Euclidean radius, within which the algorithm identifies all nearby sensor nodes and selects the one with the longest queue.
DQN: The UAV optimizes its data collection schedule using DQN by dynamically adapting to key environmental states, including: Queue length, Battery level, and Channel condition. The agent learns an optimal policy through Q-value approximation, minimizing packet loss.
DQN baseline employs a fully connected neural network with two hidden layers (400 and 300 units, ReLU activations) to map a 3-dimensional state input to a 10-dimensional action space. Training uses an $\epsilon$-greedy policy with an initial $\epsilon$ of 1.0, decaying exponentially by 0.9772 per step and clipped at 0.01. Key parameters include a learning rate of $5\times 10^{-4}$, a discount factor $\gamma$ of 0.99, a batch size of 100, and a replay buffer size of 1,000,000. A target network, updated every 3 steps, is used to stabilize learning with smooth L1 loss. The training schedule consists of 1,000 episodes, each with 30 timesteps, allowing for fast and frequent updates suitable for episodic UAV interactions.